bill533:

A 20° sector in a circle has an area of 21.5π yd². What is the area of the circle? Use 3.14 for pi. Enter your answer as a decimal in the box.

bill533:

@Hanna

Hero:

@bill533 how many degrees are there in a circle?

bill533:

360

Hero:

And what is 360/20 ?

bill533:

18

Hero:

Okay multiply that by 21.5pi

Hero:

To get the area of the circle

bill533:

is it 1215.18 or do i leave the 18 out since it asked for a decimal?

Hero:

The 18 IS a decimal

bill533:

oh ok

bill533:

What is the area of a sector with a central angle of 2π/9 radians and a radius of 16.7 ft? Use 3.14 for π and round your final answer to the nearest hundredth. Enter your answer as a decimal in the box.

bill533:

is it 97.35 square feet or 97.30 ?

Hero:

The formula for this is \(S = r\theta\)

bill533:

correct but 2*PI / 9 radians = 40 degrees

Hero:

Yes, correct, but I don't think you're supposed to convert to degrees. You're supposed to use 3.14 for pi

bill533:

oh ok

bill533:

but due to my calculations it appears the result is 30.99π ft²

Hero:

Explain how you got that

bill533:

Easy Substituting and multiplying the area

Hero:

I think I gave you the formula for something else. Sorry

bill533:

Well simply, it's like this ((2π/9)/2π) x 278.89π ft² = 30.99π ft²

Hero:

When in doubt, just use this formula for these kinds of problems: \(\dfrac{a}{360} = \dfrac{b}{\pi r^2} = \dfrac{c}{2 \pi r}\) a represents the measure of the sector b represents the area of the sector c represents the measure of the arc of the sector.

bill533:

That's what i did to calculate the area of the sector

Hero:

Koolio